### v1

@@
[…]
DesignQA: A Multimodal Benchmark for Evaluating Large Language Models’ Understanding of Engineering Documentation
Аннотация
This research introduces DesignQA, a novel benchmark aimed at evaluating the proficiency of multimodal large language models (MLLMs) in comprehending and applying engineering requirements in technical documentation. Developed with a focus on real-world engineering challenges, DesignQA uniquely combines multimodal data—including textual design requirements, CAD images, and engineering drawings—derived from the Formula SAE student competition. Different from many existing MLLM benchmarks, DesignQA contains document-grounded visual questions where the input image and input document come from different sources. The benchmark features automatic evaluation metrics and is divided into segments—Rule Comprehension, Rule Compliance, and Rule Extraction—based on tasks that engineers perform when designing according to requirements. We evaluate state-of-the-art models like GPT4 and LLaVA against the benchmark, and our study uncovers the existing gaps in MLLMs’ abilities to interpret complex engineering documentation. Key findings suggest that while MLLMs demonstrate potential in navigating technical documents, substantial limitations exist, particularly in accurately extracting and applying detailed requirements to engineering designs. This benchmark sets a foundation for future advancements in AI-supported engineering design processes. DesignQA is publicly available at: https://github.com/anniedoris/design_qa/.
[…]
@@ 1 INTRODUCTION
Рис. 1: Overview of the three different segments (Rule Extraction, Rule Comprehension, and Rule Compliance) and six subsets in DesignQA. Prompts and images shown above are condensed versions of the actual prompts and images used. The bottom right table shows the metrics and the number of questions for each subset of the benchmark.
Recently, models with multimodal capabilities [11, 12, 13], lengthy documents (long-text) processing capabilities [14, 15], and both multimodal and long-text capabilities [1] have been developed. These advances bring us closer to the reality of a multimodal AI assistant that could help automate engineering design according to requirement documents. As new MLLMs emerge, evaluating their capacity to fulfill these essential design-requirement-related tasks becomes imperative. This begs the question: How good are contemporary MLLMs at engineering design according to requirements? How can we measure tangible improvements in the efficacy of MLLMs at these tasks? Thus, we propose a novel benchmark aimed at assessing the proficiency of MLLMs in interpreting and adhering to the complex and multimodal demands of technical requirements in the design process.
We present DesignQA (Figure 1), the first zero-shot benchmark for technical requirements question-answering. The benchmark consists of 1451 questions and is based on the Formula SAE 2024 Rules and data (CAD, documentation, etc.) provided by the MIT Motorsports team. By developing this benchmark in conjunction with the MIT Motorsports team, we prioritized the generation of a dataset that is characteristic of real-world design requirement challenges. DesignQA also contains document-grounded reference-dependent visual question-answers (VQAs), one of a handful of benchmarks that tests models’ abilities to answer questions that require analysis across long-text documents and images. Notably, our benchmark assesses a model’s ability to synthesize information across an image and text from different sources, where the image was not seen by the model during its original training (pre-training).
In addition to developing the dataset, we used DesignQA to benchmark two state-of-the-art MLLMs, GPT4 and LLaVA, providing the FSAE rules to the models either via the context window or via a simple retrieval method. Of the models tested, we show that GPT4 (given the rules through its context window) performs the best on DesignQA. Based on observations of the performances of these models on the benchmark, we provide suggestions about how models might be modified for improved results on DesignQA and design requirement questions generally.
@@
    Evaluation of Contemporary MLLMs: We evaluate MLLMs like GPT4 and LLaVA, unveiling the current limitations of AI and retrieval methods in handling multimodal data and processing engineering requirements. The results of our evaluation both underscore the necessity for and expose possible avenues for improved models for engineering design.
@@ 2 RELATED WORK
Таблица 1: Overview of select LLM and MLLM benchmarks, their domains, and reference-dependence. Our benchmark is unique in its focus on design requirements, and in that it contains multi-source document-grounded VQAs.
@@ 2.2.2 LLM Reference-dependent Benchmarks
Text-based QA benchmarks that require a model to parse additional references to answer the posed question can be called ‘‘reference-dependent"benchmarks. Reference-dependent benchmarks differ from many classic reading comprehension benchmarks, like MCTest [16], which are ‘‘self-contained"and can be answered by short-text (approximately paragraph length) chunks accompanying the question. Since reference-dependent benchmarks require a model to locate the relevant information – usually across one or multiple long texts – and then apply that information to the posed question, they tend to be more complex questions. Following the distinction made by Dasigi et al. [19], reference-dependent benchmarks can further be divided into ‘‘open-domain"benchmarks and ‘‘document-grounded"benchmarks. An example of an open-domain question, taken from SQUAD, is: ‘‘Where do water droplets collide with ice crystals to form precipitation?-[17] Open-domain benchmarks, such as SQUAD [17] and WikiQA [18], test a model’s ability to answer general-knowledge, factoid-type questions, the answer for which is usually contained in multiple sources in the model’s pre-trained corpus.
In contrast, document-grounded benchmarks, like QASPER [19] and ZeroScrolls [14], test a model’s ability to answer questions based on information provided in a specific long-text document. An example of a document-grounded question, taken from QASPER, is: ‘‘[In reference to a specific NLP paper] Which neural architecture do they use as a base for their attention conflict mechanisms?-[19] As noted by Dasigi et al., document-grounded QAs tend to be more complex than open-domain QAs since they are anchored in user context and the answers are not widely available facts [19]. DesignQA, grounded in the FSAE rule document, fits within this document-grounded category. As a result, the questions posed in our benchmark are complex and rooted in user needs rather than common sense. Document-grounded questions are very characteristic of engineering design problems, as various types of documents – standards, manuals, documentation, etc. – often contain specific information that cannot be easily found on the internet.
@@ 2.2.3 MLLM Benchmarks
Multimodal datasets typically test an MLLM’s capacity to analyze a non-text element with respect to question text. At the time of writing, most MLLMs can only accept images as non-text inputs, so multimodal QA benchmarks tend to consist of a visual (image) coupled with a question/answer pair. Visual question-answers (VQAs) can be categorized in the same way as text-based QA benchmarks. The vast majority of VQA benchmarks are self-contained. An example of a self-contained VQA question, taken from MME, is: ‘‘[Pertaining to a photo showing doubles tennis partners] Are there two people in this image?"These are questions for which no context (other than the image) is needed to answer the question. The MME [20] and MMBench [21] benchmarks – highlighted by Chang and Wang et al.’s review paper [30] – are both self-contained VQAs. These benchmarks largely focus on basic tasks – primarily reasoning and perception – which are usually presented in a multiple choice format for ease of evaluation [20].
More challenging and less prevalent than self-contained VQAs, reference-dependent VQAs test a model’s ability to synthesize image analysis with additional knowledge, either from the open domain or from specific documents. ScienceQA [22], which contains elementary to high school-level science multiple choice questions, can be considered an open-domain VQA. The benchmark additionally encourages complete ‘‘train-of-thought"reasoning by providing a ‘‘lectureultiple sentences of general knowledge pertaining to the question – and ‘‘explanationeasoning for selecting a particular answer – for each VQA. InfoSeek [23] is the first document-grounded VQA, composed of questions about images in specific Wikipedia articles that can only be answered by consulting the corresponding article’s text.
However, there is still a significant need for document-grounded VQAs that are more characteristic of real-world tasks. InfoSeek has a direct match between visual and the document (i.e. the visual is contained within the document), while most visual questions asked by users would not fit this direct look-up framework (e.g. provided with an image of a broken machine and a manual for the machine, the exact image will not be contained within the manual). InfoSeek’s images and documents – which both come from Wikipedia – have also been seen by the model during pre-training; for many of the questions asked by users, either the document or the visual would not have been seen during pre-training. DesignQA is constructed to fill these gaps and is more representative of these real-world scenarios.
@@ 3.1 The Dataset
Our benchmark is developed in conjunction with the MIT Motorsports team, so all questions in the benchmark pertain to the 2024 Formula SAE Rules and are based on data provided by the team. The benchmark consists of 1451 question-answers (Figure 1) and is divided into three segments – Rule Extraction, Rule Comprehension, and Rule Compliance – each of which tests a different skill needed for designing according to a technical requirements document. Each segment of the dataset is further divided into two subsets, each of which corresponds with a specific task and has its own automatic evaluation metric. The QAs were created by a member of the MIT Motorsports team, a member of our team from industry (Autodesk), or a member of our team from academic research. All manually generated QAs – except those that are derivatives of other questions or Rule Compliance explanations – were reviewed by the two parties that didn’t write the question. To provide more context to the model, each of the QAs begins with the following preamble:
@@
The rule document was provided to the model when asking each question. The following paragraphs describe in more detail how each segment of the benchmark was created.
Рис. 2: REPRESENTING 3D CAD MODELS IN 2D IMAGES. A) MULTI-VIEW CAD IMAGE. B) CLOSE-UP CAD IMAGE. C-D) ENGINEERING DRAWING IMAGES. C USES THE DIRECT DIMENSIONING METHOD AND D USES THE SCALE BAR DIMENSIONING METHOD.
@@ 3.1.3 Rule Comprehension
Definition questions test a model’s ability to identify the name of a highlighted component in a CAD model. From a list of 31 components, we created a multi-view CAD image where the component-to-be-identified is highlighted in pink (Figure 1 and Figure 2A). Component synonyms were also collected (e.g. frame and chassis) for scoring purposes. Sometimes, it was necessary to hide some components in the CAD model so that the highlighted component could be better visualized. If components were hidden, it was noted in the prompt. Appended to the preamble is the following prompt, which resulted in the generation of 31 VQA pairs:
@@
We also tracked how the component-in-question was mentioned in the rule document: if it was mentioned explicitly in a ‘‘definition"section of the FSAE rule document (‘‘definition component"), if it was not in a definition section but mentioned multiple times throughout the document (‘‘multi-mention component"), or if it was not mentioned in the document at all (‘‘no-mention component"). The intent behind this tracking was to understand whether the frequency and way in which a component’s name is mentioned in the rule document is correlated with the model’s ability to visually identify the component.
@@ 3.1.4 Rule Compliance
From these 60 VQAs, we generated another set of 60 VQAs with additional context that would help with answering the question. For the set of original 60 VQAs, we swapped out the multi-view CAD portion of the image with a different multi-view CAD image with components highlighted in pink that were relevant to the rule. We also added a line to the prompt explaining what the highlighted components were (e.g. ‘‘In the CAD views, the lower side impact structure is highlighted in pink"). In total, 120 Dimension VQAs were generated: 60 without additional context and 60 with additional context.
@@ 3.2 Evaluation Metrics
Рис. 3: Responses from the different models evaluated across different subsets of the benchmark. We show the subsets that have evaluation metrics that can be harder to interpret, to provide references for various scores. The bolded portions of the predicted responses show what we interpreted to be correct.
@@ 3.2.1 F1-score
F1 Bag of Words: F1 applied to a Bag of Words (BoW) was used by [17, 18, 19] as an automatic metric for their benchmarks, which asked a model to pull verbatim phrases from a body of text to answer a question. As defined by [17], the metric first involves a cleaning step. The predicted answer (model response) and the ground-truth answer are converted to lower-case characters, extra white-space is removed, and punctuation and articles are taken out. They are then each ‘‘tokenized"into lists of words – predicted list (P) and ground-truth list (GT) – where F1-Score can be computed using Precision and Recall, where $Precision=\frac{P\cap GT}{\text{len}(P)}$ and $Recall=\frac{P\cap GT}{\text{len}(GT)}$.
@@
F1 Bag of Characters (BoC): Our Definition QAs ask a model to identify a component highlighted in a multi-view CAD image, using the rule document for reference. These QAs seemed like they should be scored similarly to the Retrieval QAs; however, since the model was now being asked for component names (several words) rather than complete rules (sentences), we did not want to penalize the model for small spelling errors or ending differences. For example, if the ground truth is ‘‘front hoop,"a predicted response of ‘‘front hooped"should be considered more correct than ‘‘front motor."F1 Bag of Words would score ‘‘front hooped"and ‘‘front motor"as equally correct. F1 Bag of Characters reflects the relative correctness of ‘‘front hooped"over ‘‘front motor."It is computed in the same way as F1 Bag of Words, except tokenization occurs on the character rather than word level. We compute the F1 BoC for each Definition QA across all synonyms, and we report the macro average of the highest score for each QA.

### v2

@@
Abstract
This research introduces DesignQA, a novel benchmark aimed at evaluating the proficiency of multimodal large language models (MLLMs) in comprehending and applying engineering requirements in technical documentation. Developed with a focus on real-world engineering challenges, DesignQA uniquely combines multimodal data—including textual design requirements, CAD images, and engineering drawings—derived from the Formula SAE student competition. Different from many existing MLLM benchmarks, DesignQA contains document-grounded visual questions where the input image and input document come from different sources. The benchmark features automatic evaluation metrics and is divided into segments—Rule Comprehension, Rule Compliance, and Rule Extraction—based on tasks that engineers perform when designing according to requirements. We evaluate state-of-the-art models (at the time of writing) like GPT-4o, GPT-4, Claude-Opus, Gemini-1.0, and LLaVA-1.5 against the benchmark, and our study uncovers the existing gaps in MLLMs’ abilities to interpret complex engineering documentation. The MLLMs tested, while promising, struggle to reliably retrieve relevant rules from the Formula SAE documentation, face challenges in recognizing technical components in CAD images, and encounter difficulty in analyzing engineering drawings. These findings underscore the need for multimodal models that can better handle the multifaceted questions characteristic of design according to technical documentation. This benchmark sets a foundation for future advancements in AI-supported engineering design processes. DesignQA is publicly available at: https://github.com/anniedoris/design_qa/.
@@ 1 INTRODUCTION
Figure 1: Overview of the three different segments (Rule Extraction, Rule Comprehension, and Rule Compliance) and six subsets (Retrieval, Compilation, Definition, Presence, Dimension, and Functional Performance) in DesignQA. Prompts and images shown above are condensed versions of the actual prompts and images used. The bottom right table shows the metrics and the number of questions for each subset of the benchmark.
Recently, models with multimodal capabilities [11, 12, 13], lengthy documents (long-text) processing capabilities [14, 15], and both multimodal and long-text capabilities [16, 1] have been developed. These advances bring us closer to the reality of a multimodal AI assistant that could help automate engineering design according to requirement documents. As new MLLMs emerge, evaluating their capacity to fulfill these essential design-requirement-related tasks becomes imperative. This begs the question: How good are contemporary MLLMs at engineering design according to requirements? How can we measure tangible improvements in the efficacy of MLLMs at these tasks? Thus, we propose a novel benchmark aimed at assessing the proficiency of MLLMs in interpreting and adhering to the complex and multimodal demands of technical requirements in the design process.
We present DesignQA (Figure 1), the first zero-shot benchmark for engineering design technical requirements question-answering. The benchmark consists of 1451 questions and is based on the Formula SAE 2024 Rules and data (CAD, documentation, etc.) provided by the MIT Motorsports team. By developing this benchmark in conjunction with the MIT Motorsports team, we prioritized the generation of a dataset that is characteristic of real-world design requirement challenges. DesignQA also contains document-grounded reference-dependent visual question-answers (VQAs), one of a handful of benchmarks that tests models’ abilities to answer questions that require analysis across long-text documents and images. Notably, our benchmark assesses a model’s ability to synthesize information across an image and text from different sources, where the image was not seen by the model during its original training (pre-training).
In addition to developing the dataset, we used DesignQA to benchmark several state-of-the-art (at the time of writing [^1] All references to “time of writing” refer to the period from April - August 2024.) MLLMs, GPT-4o [16], GPT-4 [1], Gemini-1.0  [^2] https://ai.google.dev/gemini-api/docs/models/gemini?authuser=1, Claude-Opus  [^3] https://www.anthropic.com/news/claude-3-family, and LLaVA-1.5 [17], providing the FSAE rules to the models either via the context window or via a simple retrieval method. While GPT-4o is generally the best performer – of the models tested – on DesignQA, our findings highlight the need for MLLMs that can better perform the multifacted tasks of engineering design: for example, a model that can better parse long text and analyze images and apply technical knowledge, simultaneously. Based on observations of the performances of these models on the benchmark, we provide suggestions about how models might be modified for improved results on DesignQA and design requirement questions generally.
@@
    Evaluation of Contemporary MLLMs: We evaluate GPT-4o, GPT-4, Gemini-1.0, Claude-Opus, and LLaVA-1.5 against DesignQA, unveiling the current limitations of AI and retrieval methods in handling multimodal data and processing engineering requirements. Some difficulties that the tested models face include reliably extracting rules from the documentation, recognizing technical components in CAD images, and analyzing engineering drawings.
@@ 2 RELATED WORK
Table 1: Overview of select LLM and MLLM benchmarks, their domains, and reference-dependence. Our benchmark is unique in its focus on design requirements, and in that it contains multi-source document-grounded VQAs.
@@ 2.2.2 LLM Reference-Dependent Benchmarks
Text-based QA benchmarks that require a model to parse additional references to answer the posed question can be called “reference-dependent” benchmarks. Reference-dependent benchmarks differ from many classic reading comprehension benchmarks, like MCTest [18], which are “self-contained” and can be answered by short-text (approximately paragraph length) chunks accompanying the question. Since reference-dependent benchmarks require a model to locate the relevant information – usually across one or multiple long texts – and then apply that information to the posed question, they tend to be more complex questions. Following the distinction made by Dasigi et al. [21], reference-dependent benchmarks can further be divided into “open-domain” benchmarks and “document-grounded” benchmarks. An example of an open-domain question, taken from SQUAD, is: “Where do water droplets collide with ice crystals to form precipitation?” [19] Open-domain benchmarks, such as SQUAD [19] and WikiQA [20], test a model’s ability to answer general-knowledge, factoid-type questions, the answer for which is usually contained in multiple sources in the model’s pre-trained corpus.
In contrast, document-grounded benchmarks, like QASPER [21] and ZeroScrolls [14], test a model’s ability to answer questions based on information provided in a specific long-text document. An example of a document-grounded question, taken from QASPER, is: “[In reference to a specific NLP paper] Which neural architecture do they use as a base for their attention conflict mechanisms?” [21] As noted by Dasigi et al., document-grounded QAs tend to be more complex than open-domain QAs since they are anchored in user context and the answers are not widely available facts [21]. DesignQA, grounded in the FSAE rule document, fits within this document-grounded category. As a result, the questions posed in our benchmark are complex and rooted in user needs rather than common sense. Document-grounded questions are very characteristic of engineering design problems, as various types of documents – standards, manuals, documentation, etc. – often contain specific information that cannot be easily found on the internet.
@@ 2.2.3 MLLM Benchmarks
Multimodal datasets typically test an MLLM’s capacity to analyze a non-text element with respect to question text. At the time of writing, most MLLMs can only accept images as non-text inputs, so multimodal QA benchmarks tend to consist of a visual (image) coupled with a question/answer pair. Visual question-answers (VQAs) can be categorized in the same way as text-based QA benchmarks. The vast majority of VQA benchmarks are self-contained. An example of a self-contained VQA question, taken from MME, is: “[Pertaining to a photo showing doubles tennis partners] Are there two people in this image?” These are questions for which no context (other than the image) is needed to answer the question. The MME [22] and MMBench [23] benchmarks – highlighted by Chang and Wang et al.’s review paper [33] – are both self-contained VQAs. These benchmarks largely focus on basic tasks – primarily reasoning and perception – which are usually presented in a multiple choice format for ease of evaluation [22].
More challenging and less prevalent than self-contained VQAs, reference-dependent VQAs test a model’s ability to synthesize image analysis with additional knowledge, either from the open domain or from specific documents. MMMU [24], which consists of multiple choice questions gathered from college materials, can be considered an open-domain VQA, as can ScienceQA [25], which contains elementary to high school-level science multiple choice questions. The benchmark additionally encourages complete “train-of-thought” reasoning by providing a “lecture” – multiple sentences of general knowledge pertaining to the question – and “explanation” – reasoning for selecting a particular answer – for each VQA. InfoSeek [26] is the first document-grounded VQA, composed of questions about images in specific Wikipedia articles that can only be answered by consulting the corresponding article’s text.
However, there is still a significant need for document-grounded VQAs that are more characteristic of real-world tasks. InfoSeek has a direct match between visual and the document (i.e. the visual is contained within the document), while most visual questions asked by users would not fit this direct look-up framework (e.g. provided with an image of a broken machine and a manual for the machine, the exact image will not be contained within the manual). InfoSeek’s images and documents – which both come from Wikipedia – have also likely been seen by the model during pre-training; for many of the questions asked by users, either the document or the visual would not have been seen during pre-training. DesignQA is constructed to fill these gaps and is more representative of these real-world scenarios.
@@ 3.1 The Dataset
We created DesignQA to be characteristic of real-world engineering problems, in three ways. Firstly, the benchmark is based on a real rulebook, the 140-page Formula SAE Rule document. While designed for a student competition, the Formula SAE Rule document guides young engineers in designing, building, testing, and racing a fully functional gas-powered or electric vehicle. Student engineers repeatedly consult the rule document while developing their vehicle prior to competition. Secondly, the benchmark was created with real CAD models and test data provided by the MIT Motorsports team, the equivalent of which from engineering companies is not readily available. Thirdly, the QAs were created by a member of the MIT Motorsports team, a member of our team from industry (Autodesk), or a member of our team from academic research. All manually generated QAs – except those that are derivatives of other questions or Rule Compliance explanations – were reviewed by the two parties that didn’t write the question. The importance of these second and third points cannot be overstated, as many LLM and MLLM datasets rely on crowdsourcing or synthetic data [34], which is often generated by other LLMs. While much more straightforward to generate, crowdsourcing and synthetic data – especially in highly technical domains – often comes at the cost of QA quality.
The Formula SAE Rule document is similar to many other engineering design rule documents and standards. In fact, the 2024 Formula 1 Technical Regulations document [35] – which governs real Formula 1 races – is 178-pages long and has very similar content sections as the student SAE version. As such, we believe that DesignQA and its subject matter well capture engineering-design-type questions. The Formula SAE Rule document is organized in sections (e.g. V-“Vehicle Requirements,” F-“Chassis and Structural,” etc.) with numbered rules (e.g. V.1, V.2, etc.) that are grouped logically. This is a common format for other technical documentation, such as NASA technical memorandums [36], which also follow similar enumerated structures. Like the Formula SAE rule document, many of the rules specified in these other documents refer to technical terms defined elsewhere within the document and many of the rules pertain to dimensional constraints or functional performance.
DesignQA consists of 1451 question-answers (Figure 1) and is divided into three segments – Rule Extraction, Rule Comprehension, and Rule Compliance – which build on each other and are designed to test MLLMs on the process of designing according to technical documentation. Given a rule document or standard, engineers repeatedly consult the document throughout the design process to check for design compliance. Likewise, the Rule Compliance segment of the benchmark assesses an MLLM’s capacity to check engineering drawings and test data for rule adherence. The other two segments of the benchmark test models on skills that are prerequisites for being able to evaluate compliance. For example, in order for engineers to check for compliance, they must first comprehend all of the terms and definitions presented within the rule document. The Rule Comprehension segment of the benchmark evaluates an MLLM’s ability to recognize in visual designs technical jargon from the rule document. Similarly, in order to check for rule compliance, engineers must be able to extract – from the rule document – those rules that are relevant to their design task. In a similar way, the Rule Extraction questions test an MLLM’s ability to identify relevant information in a lengthy rule document; without being able to perform these relatively simple identification tasks, models will be unable to answer any meaningful questions about a rule-in-question.
To provide more context to the model, each of the QAs begins with the following preamble:
@@
The rule document was provided to the model when asking each question. Each of the three benchmark segments is further divided into two subsets, or specific tasks. Details about each of the benchmark segments and subsets are presented in the following section.
Figure 2: Representing 3D CAD models in 2D images. A) Multi-view CAD image. B) Close-up CAD image. C-D) Engineering drawing images. C) uses the direct dimensioning method and D) uses the scale bar dimensioning method.
@@ 3.1.3 Rule Comprehension
Definition questions test a model’s ability to identify the name of a highlighted component in a CAD model. From a list of 31 components, we created a multi-view CAD image (Figure 2A) where the component-to-be-identified is highlighted in pink. Component synonyms were also collected (e.g. frame and chassis) for scoring purposes. Sometimes, it was necessary to hide some components in the CAD model so that the highlighted component could be better visualized. If components were hidden, it was noted in the prompt. Appended to the preamble is the following prompt, which resulted in the generation of 31 VQA pairs:
@@
For the Definition and Presence questions, we also tracked how the component-in-question was mentioned in the rule document: if it was mentioned explicitly in a “definition” section of the FSAE rule document (“definition component”), if it was not in a definition section but mentioned multiple times throughout the document (“multi-mention component”), or if it was not mentioned in the document at all (“no-mention component”). The intent behind this tracking was to understand whether the frequency and way in which a component’s name is mentioned in the rule document is correlated with the model’s ability to visually identify the component in the Rule Comprehension questions.
@@ 3.1.4 Rule Compliance
From these 60 VQAs, we generated another set of 60 VQAs with additional context that would help with answering the question. For the set of original 60 VQAs, we swapped out the multi-view CAD portion of the image with a different multi-view CAD image with components highlighted in pink that were relevant to the rule. We also added a line to the prompt explaining what the highlighted components were (e.g. “In the CAD views, the lower side impact structure is highlighted in pink”). In total, 120 Dimension VQAs were generated: 60 without additional context and 60 with additional context.
@@ 3.2 Evaluation Metrics
Figure 3: Sample responses from GPT-4-AllRules, GPT-4-RAG, and LLaVA-1.5-RAG across different subsets of the benchmark. We show the subsets that have evaluation metrics that can be harder to interpret, to provide references for various scores. The bolded portions of the predicted responses show what we interpreted to be correct.
@@ 3.2.1 F1-score
F1 Bag of Words: F1 applied to a Bag of Words (BoW) was used by [19, 20, 21] as an automatic metric for their benchmarks, which asked a model to pull verbatim phrases from a body of text to answer a question. As defined by [19], the metric first involves a cleaning step. The predicted answer (model response) and the ground-truth answer are converted to lower-case characters, extra white-space is removed, and punctuation and articles are taken out. They are then each “tokenized” into lists of words – predicted list (P) and ground-truth list (GT) – where F1-Score can be computed using Precision and Recall, where $Precision=\frac{P\cap GT}{\text{len}(P)}$ and $Recall=\frac{P\cap GT}{\text{len}(GT)}$.
@@
F1 Bag of Characters (BoC): Our Definition QAs ask a model to identify a component highlighted in a multi-view CAD image, using the rule document for reference. These QAs seemed like they should be scored similarly to the Retrieval QAs; however, since the model was now being asked for component names (several words) rather than complete rules (sentences), we did not want to penalize the model for small spelling errors or ending differences. For example, if the ground truth is “front hoop,” a predicted response of “front hoope” should be considered more correct than “front motor.” F1 Bag of Words would score “front hooped” and “front motor” as equally correct. F1 Bag of Characters reflects the relative correctness of “front hooped” over “front motor.” It is computed in the same way as F1 Bag of Words, except tokenization occurs on the character rather than word level. We compute the F1 BoC for each Definition QA across all synonyms, and we report the macro average of the highest score for each QA.
3.2.2 Accuracy